OpenStudy (jenniferjuice):

Provide a system of TWO equations in slope-intercept form with no solutions. Using complete sentences, explain why this system has no solutions.

OpenStudy (anonymous):

2 lines that never cross are?

OpenStudy (unklerhaukus):

//

OpenStudy (jenniferjuice):

yeah parallel

OpenStudy (unklerhaukus):

good, and what is the relationship between the slopes of parallel lines ?

OpenStudy (jenniferjuice):

the slope is the same ?

OpenStudy (jenniferjuice):

maybe...

OpenStudy (anonymous):

yes parallel lines share the same slope

OpenStudy (jenniferjuice):

so what would i do? how would i create an equation?

OpenStudy (unklerhaukus):

do you know the general slope-intercept form ?

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (jenniferjuice):

yeah i know general slope form @UnkleRhaukus its the y2-y1 and x2-x1 one right?

OpenStudy (anonymous):

slope intercept y=mx +b

OpenStudy (anonymous):

y2-y1 one is the point slope form

OpenStudy (jenniferjuice):

and so general slope form would be?

OpenStudy (anonymous):

i suggest rereading your question

OpenStudy (anonymous):

also reread what unkle said "general slope-intercept form"